You are purchasing pens and pencils to be sold at the school store. You need a total of 50 boxes of pens and pencils. The pens cost $3 per box and the pencils cost $2 per box. Find the total cost of the pens and pencils when you purchase 15 boxes of pens.
Answer by TimothyLamb(4379)

x = number of pen boxes = 15
y = number of pencil boxes = ?
---
x + y = 50
15 + y = 50
y = 50 - 15
y = 35
y = number of pencil boxes = 35
---
total cost = 3x + 2y
total cost = 3*15 + 2*35
---
answer:
total cost = $115
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php